janaesia:

The path traveled by a bottlenose dolphin as it jumps out of water is modeled by the equation y = −0.4x2 + 3x, where y is the height above water and x is the horizontal distance in feet. If a beam of light is shone upward at an angle modeled by the equation x + y = 10, at what height from the water's surface will the beam of light hit the dolphin?

Vocaloid:

you have to write both equations in terms of y and solve for x y = -0.4x^2 + 3x y = 10 - x -0.4x^2 + 3x = 10 - x ^ solve for x, then plug back in and find y

janaesia:

ive done it soo many times, and i just cant get it

Vocaloid:

move all the variables to one side by adding/subtracting as necessary -0.4x^2 + 3x = 10 - x subtracting 10 from both sides gives -0.4x^2 + 3x - 10 = -x adding x to both sides gives -0.4x^2 + 4x - 10 = 0 ^ apply the quadratic formula and solve for x

janaesia:

thanks